log10 x + log10 (x+21)=2
1og10[x(x+21)=2
x^2+21x=2
x^2+21x-2=0

Answer by stanbon(75887)

log10 x + log10 (x+21)=2
1og10[x(x+21)=2
That means x(x+21) = 10^2
-------
x^2+21x-100 = 0
-----
x^2+25x-4x-100 = 0
x(x+25)-4(x+25) = 0
(x+25)(x-4) = 0
Acceptable answer:
x = 4
=============================
Cheers,
Stan H.
===================